OpenStudy (anonymous):

how do i simplify 3e^loge 2 ?

zepdrix (zepdrix):

\[\huge 3e^{\log_e 2}\]Is this what it looks like?

OpenStudy (anonymous):

yes :)

zepdrix (zepdrix):

The exponential and logarithmic function are `inverses` of one another, so they essentially "undo" each other. \[\large f\left[f^{-1}(x)\right]=x\] \[\huge e^{(\log_e x)}=x\]

zepdrix (zepdrix):

So what does this give us?\[\huge e^{(\log_e 2)}=?\]

OpenStudy (anonymous):

is it 2?

zepdrix (zepdrix):

Yessss it is :) And remember we had that big ugly thing being multiplied by 3. So now our problem becomes,\[\huge 3\left(e^{\log_e 2}\right) \qquad=\qquad 3\left(2\right)\]

OpenStudy (anonymous):

ohhhhh i seee thats easy :D

OpenStudy (anonymous):

hang on ive got another question

zepdrix (zepdrix):

This idea of inverses is a little tricky. It might take a few more problems before it really sticks in your brain :)

OpenStudy (anonymous):

hahahahaa truuueee how do i simplifyy -2e^-loge2?

OpenStudy (anonymous):

i dont get how -loge2 is equal to 1/2 :(

zepdrix (zepdrix):

You'll often see the log of base e written as the "natural log". Let's do that, it'll make it a little easier to read without the e subscript. \[\huge -2e^{-\log_e2} \qquad=\qquad -2e^{-\ln2}\]

zepdrix (zepdrix):

Have you seen that before? :O If not it'll be confusing.

OpenStudy (anonymous):

i think this is sort of the first timee and i when i just saw that i just went blankk

zepdrix (zepdrix):

Ok fine fine c: we'll stick with the log notation lol

OpenStudy (anonymous):

bahahha okay

zepdrix (zepdrix):

This problem is a little trickier than the last. We're not allowed to use that same trick we did in the last problem, at least not yet. The reason is because the negative sign is in the way. We need to recall a rule of logarithms.\[\large b\cdot \log(a) \qquad=\qquad \log(a^b)\] So let's apply this rule to our exponent,\[\large -1\cdot \log(2) \qquad=\qquad ?\]

OpenStudy (anonymous):

ohhhh so its 2^-1

OpenStudy (anonymous):

rightt and thats 1/2

OpenStudy (anonymous):

ohhh i get it

zepdrix (zepdrix):

Yesss good good.

OpenStudy (anonymous):

thanks for helpinggg :)

zepdrix (zepdrix):

no probs \c:/